OpenStudy (theviper):

Please help:) The value of a & b so that \[\LARGE{\color{green}{x^3+ax^2+bx+6}}\] is divisible by (x+1) & (x-2) are??? @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

(x+1)(x-2) = x^2 - x - 2 So if x^3+ax^2+bx+6 is divisible by x^2 - x - 2 Then x^3+ax^2+bx+6 = (x+k)(x^2 - x - 2) For some value of k Expand to get... x^3+ax^2+bx+6 = (x+k)(x^2 - x - 2) x^3+ax^2+bx+6 = x(x^2 - x - 2) + k(x^2 - x - 2) x^3+ax^2+bx+6 = x^3 - x^2 - 2x + kx^2 - kx - 2k x^3+ax^2+bx+6 = x^3 - x^2+ kx^2 - 2x - kx - 2k x^3+ax^2+bx+6 = x^3 + ( -1 + k)x^2 + (-2x - k)x - 2k Do you see where to go from here?

jimthompson5910 (jim_thompson5910):

oops made a typo in the last line, it should be x^3+ax^2+bx+6 = x^3 + ( -1 + k)x^2 + (-2 - k)x - 2k

OpenStudy (theviper):

wht is k here?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Notice the constant on the left side is +6

jimthompson5910 (jim_thompson5910):

The constant on the right side is -2k

OpenStudy (theviper):

k in ur first reply:)

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (theviper):

why did u take this k & from where??

jimthompson5910 (jim_thompson5910):

oh I just introduced another variable k...well it's actually a constant

jimthompson5910 (jim_thompson5910):

Because it states that x^3+ax^2+bx+6 is divisible by x^2 - x - 2 Which means that (x^3+ax^2+bx+6)/(x^2 - x - 2) = x+k and then I rearranged terms

OpenStudy (theviper):

ok then plz solve further i m so confuse for this question

jimthompson5910 (jim_thompson5910):

The constants on both sides are equal so 6 = -2k

OpenStudy (theviper):

so wht the value of a & b??

jimthompson5910 (jim_thompson5910):

What is the value of k?

OpenStudy (theviper):

-3

jimthompson5910 (jim_thompson5910):

So from x^3+ax^2+bx+6 = x^3 + ( -1 + k)x^2 + (-2 - k)x - 2k, we see that a = -1 + k and b = -2 - k

OpenStudy (theviper):

Answer:- a=-4 b=-1

OpenStudy (theviper):

correct??

jimthompson5910 (jim_thompson5910):

a is correct, b is not

OpenStudy (theviper):

sorry 1 made a typo

jimthompson5910 (jim_thompson5910):

good a = -4 and b = 1

OpenStudy (theviper):

thank u:) have a nice day

jimthompson5910 (jim_thompson5910):

you're welcome